If a + 2b = 3c and 2a + 4b + c = 56 what is the value of c?
(A) -12
(B) -4
(C) 8
(D) 12
(E) 24

Answer by greenestamps(13198)



(1) a+2b = 3c
(2) 2a+4b+c = 56

In the second equation, "2a+4b" is "a+2b" times 2; and the first equation gives us the value of "a+2b" in terms of c:

2a+4b+c = 56
2(a+2b)+c = 56
2(3c)+c = 56
6c+c = 56
7c = 56
c = 8

ANSWER: (C) 8